1.) Joe's Burgers sells three kinds of hamburgers - Big, Giant, and the Football. The bill of materials are:

Big Burger

Giant Burger

Football Burger

1/4 lb patty

1

1/2 lb patty

1

1 lb patty

2

regular bun

1

sesame bun

1

sesame bun

1

pickle slice

1

pickle slices

2

pickle slices

4

catsup

0.1 oz

catsup

0.2 oz

lettuce

0.3 oz

onion

0.2 oz

catsup

0.2 oz

cheese

0.5 oz

onion

0.2 oz

If the product mix is 20 percent Big, 45 percent Giant, and 35 percent Footballs, and Joe sells 200 burgers per day, how much hamburger meat is used per day?

2.) Joe's Burgers sells three kinds of hamburgers - Big, Giant, and the Football. The bill of materials are:

Big Burger

Giant Burger

Football Burger

1/4 lb patty

1

1/2 lb patty

1

1 lb patty

2

regular bun

1

sesame bun

1

sesame bun

1

pickle slice

1

pickle slices

2

pickle slices

4

catsup

0.1 oz

catsup

0.2 oz

lettuce

0.3 oz

onion

0.2 oz

catsup

0.2 oz

cheese

0.5 oz

onion

0.2 oz

If the product mix is 20 percent Big, 45 percent Giant, and 35 percent Footballs, and Joe sells 200 burgers per day, how many pickle slices are needed per day?

3.) Develop an MRP record for six periods using the following information:

period

1

2

3

4

5

6

gross requirements

20

20

40

30

30

30

scheduled receipts

50

projected available balance

2

planned order receipt

planned order release

Lead time = 1 period, lot size = 50 units, safety stock = 0 units, current inventory (prior to period 1) = 2 units, scheduled receipts = 50 units in period 1.

What is the ending value of the projected available balance (in period 6)?
Answer #1

.(a) Number of each type of burgers sold

Big burger = 200x0.2 =40

Giant burger = 200x0.45 =90

Football burger = 200x0.35 =70

Hamburger consumed = 1/4 x40+1/2x90+1x70 = 125

(b)

Pickle slices used = 1x40+2x90+4x70 = 500

(c)

Period 1 2 3 4 5 6
Gross reuirements 20 20 40 30 30 30
Scheduled receipts 50 0 0 0 0 0
projected available balance 32 12 22 42 12 32
Planned order receipt 0 0 50 50 0 50
Planned order release 0 50 50 0 50 0